A particle is moving along a circular path of radius R in such a way that at any instant magnitude of radial acceleration and tangential acceleration are equal. If at t=0 velocity of particle is V0. Find the speed of the particle after time t= R/2V0.

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Radial and tangencial acceleration 
at = dv/dt 
ar = v^2 /r 
at = ar 
dv/dt = v^2 /r 
c = 1/vo 
-1/v = t/R - 1/vo 
t = R (1/vo - 1/v)